# SIGNeRF: Scene Integrated Generation for Neural Radiance Fields

### 03 Jan 2024 v1

@@ 2 Related Work
Figure 3: Object insertion and object modification. – (top) The cow geometry is centrally placed on a meadow to obtain a photorealistic model. (middle) Note how occlusions are properly handled when generating the synthetic house based on the inserted proxy. (bottom) Objects can easily be transformed based on a prompt. Due to the more complex surface texture and geometric changes the pirate and the Batman costume required an additional iteration to obtain the same level of consistency compared to the simpler sports clothes.
@@ 3.2 Controlled Consistent Generation
Figure 5: Qualitative Comparison – SIGNeRF results are compared to Instruct-NeRF2NeRF [13] (top) and DreamEditor [64] (bottom). For the bear, the generated fur texture with SIGNeRF (left) shows a more distinguished structure and the snout regions is clearly more consistent. Compared to DreamEditor the images are different but the image quality comparible.
@@ 4.2 Qualitative Evaluation
The impact or the proxy shape and geometric detail is visualized in  Fig. 6. The shape of the proxy mesh does influence the final object as ControlNet tries to fit the depth condition created with the proxy mesh, as apparent when the body of the cow is approximated by a cylinder. Nevertheless, the results indicate that the proxy mesh does not need to be a detailed representation of the object, as the diffusion process added additional geometry details. Both low and high-poly versions yield satisfactory results.
@@ 4.3 Comparison
Another aspect is the time spent in the generation process. Instruct-NeRF2NeRF and DreamEditor, require more than an hour for a generation that is only fully visible at the end. While SIGNeRF only takes half the time on a single GPU, using one dataset iteration, it additionally provides a preview option with the reference sheet, allowing the user to adjust the output until satisfied with the results before starting the generation process. Furthermore, the image set update (Step 6) can easily be parallelized over all images in our approach while the interlocked NeRF/image updates in the other methods are purely sequential.
@@ 5 Conclusion
While being developed for NeRF scenes, shifting to other representations like Gaussian Splats [21] should be straightforward due to the modular design.
Although we are aware that this work might also be misused to forge convincing 3D scenes we hope that it overall will provide another step towards democratizing the complex task of 3D content generation.
@@ Acknowledgements
This work has been partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC number 2064/1 - project number 390727645 and SFB 1233 - project number 276693517. It was supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A and Cyber Valley
@@ Appendix A Implementation Details
Appendix B Viewer
Our development includes a NeRF viewer designed to examine the trained NeRF scenes and facilitate the placement of proxy objects or the selection of scene elements for modification. The viewer’s backend is based on Nerfstudio [50], while the frontend is entirely novel, incorporating a suite of advanced features. A depiction of the viewer interface is presented in Fig. B9. Constructed as a React [29] application, it leverages React Three Fiber [15], a React renderer for Three.js [5], to facilitate 3D scene manipulation. The Nerfstudio backend streams the NeRF data to the frontend, where it is displayed centrally in the viewport. We have augmented the backend to transmit not just the scene render but also the corresponding depth map. This data enables us to utilize a shader that merges the NeRF render with Three.js’s native rendering, allowing for the visualization of occluded elements within the NeRF scene in real-time.
Figure B9: Viewer – Outlining the placement of a proxy object within an existing NeRF scene. The viewer is divided into a 3D viewport at the center, a scene and selection column on the left and a selection specific control column on the right.
@@ Appendix B Viewer
Appendix C Dataset Generation
Figure C10: Dataset Generation – Illustrating the consistency within edited views of an original NeRF dataset. The first row shows images from the reference sheet, the rest are the views generated with the reference sheet.
SIGNeRF generates an updated NeRF image dataset with the provided reference sheet to edit a NeRF scene. Our generative pipeline (Fig. 2) demonstrates this with a selection of modified dataset images, but a completely revised dataset is not showcased. In Fig. C10, we exhibit a fully edited dataset of the ‘field’ scene (Fig. 3). Consistency is maintained across the generated images; however, some outliers might occur. These inconsistencies are minor over the dataset and are generally resolved by the NeRF optimization process, which learns to ignore such outliers. For complex scenes, the use of an LPIPS [62] loss is beneficial to further improve the consistency during the optimization.

### 27 Mar 2024 v2

@@ 2 Related Work
Figure 3: Object insertion and object modification. – (top) The cow geometry is centrally placed on a meadow to obtain a photorealistic scene. (middle) Note how occlusions are properly handled when generating the synthetic house based on the inserted proxy. (bottom) Objects can easily be transformed based on a prompt. Due to the more complex surface texture and geometric changes the pirate and the Batman costume required an additional iteration to obtain the same level of consistency compared to the simpler sports clothes.
@@ 3.2 Controlled Consistent Generation
Figure 5: Qualitative Comparison – SIGNeRF results are compared to Instruct-NeRF2NeRF [14] (top) and DreamEditor [66] (bottom). For the bear, the generated fur texture with SIGNeRF (left) shows a more distinguished structure and the snout regions is clearly more consistent. Compared to DreamEditor the images are different but the image quality comparable.
@@ 4.2 Qualitative Evaluation
The impact of the proxy shape and geometric detail is visualized in  Fig. 6. The shape of the proxy mesh does influence the final object as ControlNet tries to fit the depth condition created with the proxy mesh, as apparent when the body of the cow is approximated by a cylinder. Nevertheless, the results indicate that the proxy mesh does not need to be a detailed representation of the object, as the diffusion process added additional geometry details. Both low and high-poly versions yield satisfactory results.
@@ 4.3 Comparison
Another aspect is the time spent in the generation process. Instruct-NeRF2NeRF and DreamEditor, require more than an hour for a generation that is only fully visible at the end. While SIGNeRF only takes half the time on a single GPU, using one dataset iteration, it additionally provides a preview option with the reference sheet, allowing the user to adjust the output until satisfied before starting the generation process. Furthermore, the image set update (Step 6) can easily be parallelized over all images in our approach while the interlocked NeRF/image updates in the other methods are purely sequential.
@@ Acknowledgements
Funded by EXC number 2064/1 – Project number 390727645. This work was supported by the German Research Foundation (DFG): SFB 1233, Robust Vision: Inference Principles and Neural Mechanisms, TP 02, project number: 276693517. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Jan-Niklas Dihlmann.
@@ Appendix A Implementation Details
Appendix B Multiview Consistency
Figure B9: Inpainting Ablation & LoFTR Features – grid inpainting generation capabilities with & w/o ControlNet under certain denoising levels (label top right). LoFTR keypoint correspondences reflect the consistency between the sub images (last row). Zoom in for details.
We want to clarify that ControlNet [63] was not fine-tuned in any way. We found the multi-view consistency property empirically by exploration. Initially, combining two images, and later discovering how to reach greater consistency using a grid. The reason why diffusion models show this multi-view consistency isn’t entirely clear, but we have some hypotheses. The grid-based inpainting can achieve a degree of multi-view consistency without ControlNet, although the results become less consistent at higher denoising levels (Fig. B9), therefore only allowing subtle edits. This suggests that StableDiffusion [41] inherently supports some level of multi-view consistency and 3D understanding as also shown by [8, 62]. We assume that attention plays a crucial role, further enforced by the inpainting mechanism focusing on the relevant parts and ignoring scene elements by masking. The depth condition in ControlNet further acts as a reference for the model to infer orientation/shape/scale, even for high denoising levels as the condition is not noised. Fine-tuning to achieve even better results is interesting future work, but we think that one of the most attractive properties of our approach is the possibility of achieving high-quality generation with off-the-shelf image diffusion models.
Further there is no direct measurement of the quality of the multiview consistency within the reference sheet. This is also nontrivial to measure, else we could optimize for it in an self supervised manner. We experimented with LoFTR [50] correspondences between original, reference-sheet generated and independent generated images. Averaged results show that independent generation has $70\%$ less correspondences than the original images, while the reference sheet generated images achieve $15\%$ less correspondences. Nevertheless, we believe the supplementary videos we provided substantiate our claims of achieving multiview consistency. The feasibility of conducting 3D reconstructions from the generated images serves as a testament to this consistency. It is important to acknowledge that while the images may not exhibit perfect consistency, the underlying NeRF counteracts that by aligning 3D consistent regions and discarding outliers.
Appendix C Viewer
Our development includes a NeRF viewer designed to examine the trained NeRF scenes and facilitate the placement of proxy objects or the selection of scene elements for modification. The viewer’s backend is based on Nerfstudio [51], while the frontend is entirely novel, incorporating a suite of advanced features. A depiction of the viewer interface is presented in Fig. C10. Constructed as a React [29] application, it leverages React Three Fiber [16], a React renderer for Three.js [5], to facilitate 3D scene manipulation. The Nerfstudio backend streams the NeRF data to the frontend, where it is displayed centrally in the viewport. We have augmented the backend to transmit not just the scene render but also the corresponding depth map. This data enables us to utilize a shader that merges the NeRF render with Three.js’s native rendering, allowing for the visualization of occluded elements within the NeRF scene in real-time.
Figure C10: Viewer – Outlining the placement of a proxy object within an existing NeRF scene. The viewer is divided into a 3D viewport at the center, a scene and selection column on the left and a selection specific control column on the right.
@@ Appendix C Viewer
Appendix D Dataset Generation
Figure D11: Dataset Generation – Illustrating the consistency within edited views of an original NeRF dataset. The first row shows images from the reference sheet, the rest are the views generated with the reference sheet.
SIGNeRF generates an updated NeRF image dataset with the provided reference sheet to edit a NeRF scene. Our generative pipeline (Fig. 2) demonstrates this with a selection of modified dataset images, but a completely revised dataset is not showcased. In Fig. D11, we exhibit a fully edited dataset of the ‘field’ scene (Fig. 3). Consistency is maintained across the generated images; however, some outliers might occur. These inconsistencies are minor over the dataset and are generally resolved by the NeRF optimization process, which learns to ignore such outliers. For complex scenes, the use of an LPIPS [64] loss is beneficial to further improve the consistency during the optimization.